# Approximating the order 2 quantum Wasserstein distance using the moment-SOS hierarchy

### v1

@@ 3.2 A hierarchy of moment-SOS relaxations
We now approximate the problem defined in (2) using finite-dimensional semidefinite programs. We consider a hierarchy of moment-SOS relaxations of the GMP instance (2). The hierarchy of moment relaxations, indexed by $t\geq 2$, is
$W^{2}_{2}(\rho,\sigma)_{t}:=\inf_{L}$ $L(f)$ (6)
@@
For $t=\infty$, the parameter $W_{2}^{2}(\rho,\nu)_{\infty}$ involves linear functionals acting on the full polynomial space $\mathbb{C}[\mathbf{x},\overline{\mathbf{x}},\mathbf{y},\overline{\mathbf{y}}]$. Note that for all integer $t\geq 2$, the moment program (6) is a relaxation of (2), thus $W_{2}^{2}(\rho,\nu)_{t}\leq W_{2}^{2}(\rho,\nu)$. In addition, one has $W_{2}^{2}(\rho,\nu)_{t}\geq D(\rho,\nu)_{t}$ by weak duality. The next result states that strong duality holds and that the moment-SOS hierarchy converges to the order 2 quantum Wasserstein distance.
@@
To prove that $\mathcal{R}_{t}^{\operatorname{opt}}$ is bounded, we rely on the first result of Lemma 1. Let $L_{t}\in\mathcal{R}_{t}$. Note that $L_{t}(1)=L_{t}(\operatorname{Tr}(\mathbf{x}\mathbf{x}^{*}))=\operatorname{Tr}(\rho)=1$. The assumptions of Lemma 1 are satisfied since $2-\|\mathbf{x}\|-\|\mathbf{y}\|^{2}\in\mathcal{Q}(\mathbf{X})_{2}$ and $L_{t}$ is nonnegative on $\mathcal{Q}(\mathbf{X})_{2t}$. Then we have $|L_{t}(w)|\leq 2^{|w|/2}$ for all $w\in[\mathbf{x},\overline{\mathbf{x}},\mathbf{y},\overline{\mathbf{y}}]_{2t}$. Thus the sets $\mathcal{R}_{t}$ and $\mathcal{R}_{t}^{\operatorname{opt}}$ are both bounded.
@@
Remark 11.
Similarly to Remark 9, the dual SOS program (7) is not attained. Indeed, for any feasible solution $(\Lambda,\Gamma)$ of (7), the tuple $(\Lambda+\alpha I,\Gamma-\alpha I)$ is also feasible for any arbitrary large $\alpha\in\mathbb{R}$ since $\alpha(\|x\|^{2}-\|y\|^{2})\in\mathcal{Q}(X)_{2t}$ for any $t\in\mathbb{N}$.
@@ 4 Concluding numerical experiments and perspectives
A first topic of further investigation would be to reduce the computational effort required to solve the pair of semidefinite programs (6)-(7) at a fixed relaxation order $t$. For this, one could rely on the alternative formulations proposed in [GLS22, Section 5.2] that allow us to block-diagonalize the matrix variable in various ways; see in particular [GLS22, Eq. (61) and Lemma 22].
@@
This note focused solely on the order $p=2$ quantum Wasserstein distance. An exciting research track would be to identify for which other values of $p$ one can characterize the corresponding quantum Wasserstein distances as (non)linear GMP instances.
@@ Acknowledgments.
We thank Emily Beatty, Jonas Britz and Monique Laurent for fruitful discussions. Part of this research was performed while the first author was visiting the Institute for Pure and Applied Mathematics (IPAM), which is supported by the National Science Foundation (Grant No. DMS-1925919). This work was also supported by the European Union’s HORIZON–MSCA-2023-DN-JD programme under the Horizon Europe (HORIZON) Marie Sklodowska-Curie Actions, grant agreement 101120296 (TENORS).

### v2

@@ 3.2 A hierarchy of moment-SOS relaxations
We now approximate the problem defined in (2) using finite-dimensional semidefinite programs. We consider a hierarchy of moment-SOS relaxations of the GMP instance (2). Note that the minimal relaxation order should be $t=2$ since the objective function has degree $4$. The hierarchy of moment relaxations, indexed by $t\geq 2$, is
$W^{2}_{2}(\rho,\sigma)_{t}=\inf_{L}$ $L(f)$ (6)
@@
For $t=\infty$, $W_{2}^{2}(\rho,\nu)_{\infty}$ involves linear functionals acting on the full polynomial space $\mathbb{C}[\mathbf{x},\overline{\mathbf{x}},\mathbf{y},\overline{\mathbf{y}}]$. Note that for all integer $t\geq 2$, the moment program (6) is a relaxation of (2), thus $W_{2}^{2}(\rho,\nu)_{t}\leq W_{2}^{2}(\rho,\nu)$. In addition, one has $W_{2}^{2}(\rho,\nu)_{t}\geq D(\rho,\nu)_{t}$ by weak duality. The next result states that strong duality holds and that the moment-SOS hierarchy converges to the order 2 quantum Wasserstein distance.
@@
To prove that $\mathcal{R}_{t}^{\operatorname{opt}}$ is bounded, we rely on the first result of Lemma 1. Let $L_{t}\in\mathcal{R}_{t}$. Note that $L_{t}(1)=L_{t}(\operatorname{Tr}(\mathbf{x}\mathbf{x}^{*}))=\operatorname{Tr}(\rho)=1$. The assumptions of Lemma 1 are satisfied since $2-{\color[rgb]{0,0,0}\|\mathbf{x}\|^{2}}-\|\mathbf{y}\|^{2}\in\mathcal{Q}(\mathbf{X})_{2}$ and $L_{t}$ is nonnegative on $\mathcal{Q}(\mathbf{X})_{2t}$. Then we have $|L_{t}(w)|\leq 2^{|w|/2}$ for all $w\in[\mathbf{x},\overline{\mathbf{x}},\mathbf{y},\overline{\mathbf{y}}]_{2t}$. Thus the sets $\mathcal{R}_{t}$ and $\mathcal{R}_{t}^{\operatorname{opt}}$ are both bounded.
@@ 4 Concluding numerical experiments and perspectives
A first topic of further investigation would be to reduce the computational effort required to solve the pair of semidefinite programs (6)-(7) at a fixed relaxation order $t$. For this, one could rely on the alternative formulations proposed in [GLS22, Section 5.2] that allow us to block-diagonalize the matrix variable in various ways; see in particular [GLS22, Eq. (61) and Lemma 22]. In addition to this computational effort reduction, another practical goal is the extraction of optimal quantum transport plans. When a pseudo-moment matrix obtained at a given relaxation order satisfies the so-called flat extension condition, one can conclude that the corresponding hierarchy has finite convergence. Under this condition, the algorithm from [HL05] outputs the atomic decomposition of the optimal measure. While this condition did not hold in our numerical experiments, one could derive another adapted one to moment relaxations with block-diagonalized matrix variables.
@@
This note focused solely on the order $p=2$ quantum Wasserstein distance. Other cases of special interest include the infinite-order quantum Wasserstein distance ($p=\infty$). An exciting research track would be to identify for which values of $p$ one could characterize the corresponding quantum Wasserstein distances as linear or more generally convex moment problems. Additionally, the recent framework [KMV25] would allow us to also consider quantum transport problems cast as nonlinear polynomial moment problems.
@@ Acknowledgments.
The authors are grateful to the two anonymous referees for their constructive feedback. We thank Emily Beatty, Jonas Britz and Monique Laurent for fruitful discussions. Part of this research was performed while the first author was visiting the Institute for Pure and Applied Mathematics (IPAM), which is supported by the National Science Foundation (Grant No. DMS-1925919). This work was also supported by the European Union’s HORIZON–MSCA-2023-DN-JD programme under the Horizon Europe (HORIZON) Marie Sklodowska-Curie Actions, grant agreement 101120296 (TENORS).
Data availability statement.
We do not analyse or generate any datasets, because our work proceeds within a theoretical and mathematical approach.